DIRC Operations Weekly Report

Period: Saturday 2006/07/15 - Friday 2006/07/21

Oncall: Jerome Bequilleux
Fast Monitoring: Bradley Wogsland, Jerome Bequilleux

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060715.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060715 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 11.2*1033 (previous week: 10.2*1033)

BaBar integrated lumi: 3.13 fb-1 (previous week: 3.98 fb-1)

Run 5b PEP delivery: 62.49 fb-1

Run 5 total delivery: 132.63 fb-1 Run 5 total BaBar recorded: 127.93 fb-1 including 9.82 fb-1 off-peak data

PEP activities this week:

  • All the week, PEPII tried to find the reason of HER aborts (appeared in April but became significant in the middle of May).
       -tried to change the bunch pattern (keep aborts at 1200mA; vacuum activity) -> back to normal bunch pattern
       -tried to raise HER current with increasing SVT RAD threshold (BABAR ramped down).
       -change of vacuum cooling system
       -studied correlations between belows' temperature and aborts.
  • Thursday 13th: Leak of compressed air outside the gas shack (Blg 636).
  • Friday 14th: UPS alarms -> an expert will come to have a look and hopelly fix the problem.

    PEP activities planned next week :

    Foreseen work opportunities (PEP down periods):

    BaBar activities/issues this week:

    DIRC activities

    See DIRC logbook for details

  • Degasing pump (following)
           -Monday 10th: Jerry checked the situation with the degassing pump (we had glitches 5 times over the week-end). We will continue like that, knowing that the 52V readouts for the degassing pump in ambient is not a hardware problem of the degassing pump. As we keep a lower alarm for this variable, we will still know if the degassing pump is having troubles.

  • Missing documentation:
           -Tuesday 11th: talking with Philippe(Grenier) about the DIRC analogic chips, for data acquisition, we found out that corresponding thresholds need to be set when changing a DFB. Apparently, no documentation exists about it...

  • UPS work:
           - Friday 14th: -Matt moved the degasing pump back to the UPS system (was put on normal power when there were problems on the UPS) at 8am. Maybe it will partially fix the glitches problem with the readout of the degassing pump variable. However, as ths UPS is having trouble again, and will need some work tomorrow, Matts switch back to normal power at 3 pm.

  • Fan Monitoring (following):
           - Friday 14th: Nicolas changed daily fan monitoring plots. We keep the usual scale, but there will be arrows for values that go above or under this scale (see Fans Monitoring Plots ).

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Manpower situation

    Links